OpenStudy (anonymous):

A square and a circle both made out of wire enclose an area of 45.35cm^2. Find when the wires are flattened out, which wire is longer and by how much? Round to the nearest hundredths.

OpenStudy (anonymous):

use the equation of the circle to find the radius,then multiply the radius by 2. that will be the diameter

OpenStudy (binary3i):

\[A=pi R^2 = L^2\] R radius of circle and L length of a side of squre.\[C=pi 2R = 2\sqrt{pi A} \] \[4L= 4\sqrt{A}\]

OpenStudy (anonymous):

I am a little confused.

OpenStudy (binary3i):

4L is the length of square and C is the length of circle, of wires actually.

OpenStudy (anonymous):

I mean to find the radius?

OpenStudy (binary3i):

A is 45.35cm^2 pi is 3.141592

OpenStudy (anonymous):

Can you help me with the rest of this I am so confused?

OpenStudy (binary3i):

R is the radius L is the side length\[45.35cm^2 = 3.14 \times R^2\] \[\sqrt{45.35/3.14}= R\] now the length of the circle is C\[C= 2\times 3.14\times R\] \[C=2 \times 3.14 \times \sqrt{45.35/3.14}= 2 \times \sqrt{3.14 \times 45.35}\] now the square \[45.35= L^2\] \[L=\sqrt{45.35}\] length of the square is 4L\[length of the \square = 4L= 4 \sqrt{45.35}\] and now you compair C and 4L

OpenStudy (anonymous):

I see it now, thanks.